What is the concentration of 10.0 mL of HBr if it takes 5.0mL of a 0.255M LiOH solution to neutralize it?
Answer #1

Liot & HR8 wirrt 2 HBO (1) - 10.0m2 - 10.04 0.01L 10004 con of HBV (0) - M concentration of wott = 0.2554 - M. veure of not =